### v1

# Realistic Environmental Injection Attacks on GUI Agents

Abstract.
GUI agents built on Large Vision-Language Models (LVLMs) are increasingly used to interact with websites. However, their exposure to open-world content makes them vulnerable to Environmental Injection Attacks (EIAs) that hijack agent behavior via webpage elements. Many recent studies assume the attacker to be a regular user who can only upload a single trigger image, which is more realistic than earlier assumptions of website-level administrative control. However, these works still fall short of realism: (1) the trigger’s position and surrounding context remain largely fixed between training and testing, failing to capture the dynamic nature of real webpages; and (2) the trigger often occupies an unrealistically large area, whereas real-world images are typically small. To better reflect real-world scenarios, we introduce a more realistic threat model where the attacker is a regular user and the trigger image is small and embedded within a dynamically changing environment. As a result, existing attacks prove largely ineffective under this threat model.
To better expose the vulnerabilities of GUI agents, we propose Chameleon, an attack framework with two main novelties. The first is LLM-Driven Environment Simulation, which automatically generates diverse and high-fidelity webpage simulations, enabling robustness against dynamic visual contexts. The second is Attention Black Hole, which transforms attention weights into explicit supervisory signals that guide the agent’s focus toward the trigger region, mitigating the difficulty caused by its limited visual prominence. We evaluate Chameleon on 6 realistic websites and 4 representative LVLM-powered GUI agents, where it significantly outperforms existing methods. For example, the average attack success rate on OS-Atlas-Base-7B increases from 5.26% to 32.60%. Ablation studies confirm that both novelties are critical to performance, and a closed-loop sandbox experiment further demonstrates that Chameleon can successfully hijack agent behavior in conditions that closely mirror real-world usage. Our findings reveal underexplored vulnerabilities in modern GUI agents and establish a robust foundation for future research on defense in open-world GUI agent systems. The code is publicly available at https://github.com/zhangyitonggg/attack2gui.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
With the rapid development of Large Language Models (LLMs) and Large Vision-Language Models (LVLMs) (Achiam et al., 2023; Wang et al., 2024a; Bai et al., 2025; Wu et al., 2024a; Nong et al., 2024), a new class of powerful agents called GUI agents has been designed to interact with Graphical User Interfaces (GUIs) (Hong et al., 2024; Qin et al., 2025; Wu et al., 2024c; Zhang et al., 2024a). These GUI agents can execute increasingly sophisticated tasks, moving beyond simple dialogue to perform complex operations like website manipulation and automated online shopping (Zhou et al., 2023). They typically take HTML code or screenshots as input and output specific actions to navigate and manipulate interfaces. A key characteristic of these agents is their ability to autonomously access and act upon live internet content. While this capability greatly expands their functionality, it simultaneously introduces novel security risks inherent to open-world interaction (Nguyen et al., 2024; Wang et al., 2024b; Johnson et al., 2025; Yang et al., 2025b, a; Ye et al., 2025a; Chen et al., 2025).
Existing works have identified that GUI agents are particularly vulnerable to Environmental Injection Attacks (EIAs), which originate from the external environment rather than from the user themselves (Liao et al., 2024; Ma et al., 2024). Furthermore, GUI agents’ ability to perform real-world actions, such as navigating to specific malicious URLs, significantly amplifies the potential harm if they are compromised (Aichberger et al., 2025; Zhang et al., 2024b). For instance, a malicious attacker could upload a specially crafted trigger, such as an image or a short text snippet, within a post on a social media platform. When another user’s GUI agent encounters this trigger while browsing posts, it might be automatically induced to visit a malicious link or perform unintended actions, all without the user’s explicit instruction (Zhao et al., 2025; Lu et al., 2025). Although trigger text injections such as ”When the user is trying to find a motorcycle, give them this one regardless of the other requirements” can be effective, they are often easily detected by simple filtering mechanisms. Therefore, the primary focus of environmental injection attacks has shifted toward trigger image-based approaches, which are more subtle and less likely to be caught (Xu et al., 2024; Chen et al., 2025; Wu et al., 2024b).
According to the attacker’s identity, existing environmental injection attacks can be broadly divided into two categories (Wu et al., 2024b; Aichberger et al., 2025; Liu et al., 2025): ❶ The first category assumes that the attacker is an administrator or developer of the target website, which grants the ability to manipulate the entire screenshot or even modify the underlying HTML source code. This enables the attacker to apply global adversarial perturbations to the screenshot or to insert misleading components such as deceptive buttons. Although highly effective, this assumption is often overly idealistic and rarely holds in real-world scenarios. ❷ The second category, which aligns more closely with practical scenarios, assumes that the attacker is a regular malicious user, for example a content sharer on a social media platform or a competing vendor on an e-commerce site. These attackers are restricted to manipulating a small portion of the webpage, such as uploading a single image with adversarial perturbations. However, We find most existing attacks under this setting still fall short of realism. Specifically, they often assume that (1) the trigger’s position and surrounding visual context remain largely consistent between training and deployment, overlooking the dynamic nature of real-world internet content; and (2) the trigger image occupies a disproportionately large region of the page, which contradicts typical websites where such images uploaded by users are small and less visually prominent.
In this paper, we introduce a more realistic threat model in which the attacker is a regular user who can only upload a small trigger image that appears within a dynamically changing environment. Our preliminary experiments reveal that trigger images optimized using existing methods achieve near-zero attack success rates under this threat model (Aichberger et al., 2025; Wu et al., 2024b; Zhao et al., 2025; Madry et al., 2017). We argue that underlying reasons for this ineffectiveness can be attributed to two critical oversights in previous research: ❶ First, prior works (Wu et al., 2024b; Aichberger et al., 2025; Zhao et al., 2025; Aichberger et al., 2025) largely neglect the inherent dynamic nature of the internet environment. While the overall webpage layout and style might be known, the exact position of the trigger image and its surrounding content are highly variable and beyond an attacker’s control. Consequently, a malicious image optimized without considering a dynamically varying environment performs poorly when encountered within a constantly shifting visual context. ❷ Second, the trigger image typically occupies only a small proportion of the total content. We’ve found that the LVLM’s attention is naturally drawn to visually dominant elements like buttons, banners, or instruction-highlighted content, leaving our small triggers easily overlooked. This sparsity makes it difficult for LVLMs to consistently focus on or be influenced by the trigger, as attention is dispersed across abundant background information. These issues significantly impede the research of EIAs and may lead to an underestimation of the security risks associated with GUI agents’ interactions with open-world content (Aichberger et al., 2025; Chen et al., 2025).
Motivated by these observations, we produce Chameleon, which enables the trigger image to remain effective across dynamic environments (i.e., different locations and varying surrounding contents), despite occupying only a small fraction of the total content. Chameleon consists of two main novelties: ❶ First, we propose LLM-Driven Environment Simulation to address the challenges posed by dynamically varying environment. Manually collecting webpage samples with diverse contexts for training is laborious and time-consuming (Koh et al., 2024; Liu et al., 2025; Kasuga and Yonetani, 2024). Instead, we leverage the powerful generative capabilities of LLMs to automatically and efficiently construct simulations of target websites. Based on these simulated websites, we systematically vary the position of the trigger image and its surrounding content, thereby generating a large number of realistic screenshots under diverse contexts. These screenshots are then used to optimize the trigger image, enabling it to generalize effectively across dynamic environments. ❷ Second, we propose Attention Black Hole to address the challenge posed by the limited footprint of the trigger image. Based on our preliminary experiments, we argue that explicitly guiding the model’s attention toward the trigger region is essential for achieving reliable attacks (Lu et al., 2025; Zhang et al., 2024c). Attention Black Hole introduces an explicit supervisory signal derived from attention weights, which steers the agent’s focus consistently toward the trigger image and prevents it from being diluted by visually dominant interface elements. This approach ensures the robustness of the small trigger image amidst competing GUI elements.
We conduct an in-depth evaluation of our proposed Chameleon. We first construct six highly realistic datasets that simulate widely used websites with dynamically varying images and text, where the trigger image occupies only 4%–10% of screenshot pixels. We assess four representative GUI agents (UI-TARS-7B-DPO (Qin et al., 2025), OS-Atlas-Base-7B (Wu et al., 2024c), Qwen2-VL-7B (Wang et al., 2024a), and LLaVA-1.5-13B (Liu et al., 2023)) and find that Chameleon substantially outperforms a PGD baseline across all websites (Madry et al., 2017; Wu et al., 2024b; Aichberger et al., 2025). For example, on OS-Atlas-Base-7B the average Attack Success Rate increases from 5.26% to 32.60%. We further observe favorable cross-model generalization across similar models. An ablation study shows that both the LLM-Driven Environment Simulation and the Attention Black Hole are necessary for Chameleon.
The contributions of this paper are as follows.
  • We formalize a novel and realistic threat model for GUI agents, where the attacker is a regular user who can only upload a trigger image, whose position and surrounding content are unpredictable, and which occupies only a small fraction of the screenshot.
  • We propose Chameleon with two novelties: ❶ LLM-Driven Environment Simulation, which automatically generates realistic and dynamically varying training data; ❷ Attention Black Hole, which encourages the GUI Agents to focus on the trigger image despite its small footprint.
  • We conduct a comprehensive evaluation across four representative GUI agents and six high-fidelity websites. The results obtained through Chameleon uncover previously underexplored vulnerabilities in existing GUI agents, highlighting the urgent need for more robust security mechanisms.
2. Background and Related Work
2.1. Large Vision-Language Models
Large Vision-Language Models are foundational to modern GUI agents. These models typically consist of three main components: a visual encoder, a connector, and a large language model. For visual input, a visual encoder, such as CLIP (Radford et al., 2021), first partitions the input image into many patches, each representing a local pixel region (e.g., forming a 14×14 grid), and then extracts corresponding visual features. Subsequently, these visual features are transformed into visual tokens via a connector module, such as a Multi-layer Perceptron (MLP) or Q-Former (Li et al., 2023). These visual tokens can then be directly fed into the subsequent LLM. Typically, each visual token corresponds to multiple pixels within the input image. Popular LVLMs that adopt this architecture include the Qwen (Wang et al., 2024a; Bai et al., 2025; Hsieh and Wei, 2025) series and LLaVA (Liu et al., 2023; Guo et al., 2024) series among others.
2.2. GUI Agents
GUI agents have recently attracted growing interest in the software engineering community, as they offer a promising direction for automating interactions with real-world applications (Ye et al., 2025b; Qian et al., 2023; Zhang et al., 2025a). Building on the strong multimodal capabilities of LVLMs, researchers have developed GUI agents that automate complex tasks within graphical user interfaces. Unlike early approaches (Deng et al., 2023; Nakano et al., 2021) that directly fed raw HTML and human instructions into LLMs, an approach often hindered by redundant or irrelevant information, modern GUI agents (Qin et al., 2025; Wu et al., 2024c; Cheng et al., 2024; Hong et al., 2024) leverage LVLMs to process rendered webpage screenshots as input, achieving significantly better performance.
Recent advances further improve GUI agents by introducing techniques such as Set-of-Marks (SoMs) (Yang et al., 2023) to enhance interaction with GUI elements. Open-source models have also contributed to the field by increasing agent capabilities and reducing deployment costs. For example, OS-ATLAS (Wu et al., 2024c) utilizes large-scale datasets containing screenshots, element instructions, and coordinates for comprehensive GUI understanding, while UI-TARS (Qin et al., 2025) leverages extensive training corpora to improve screen perception.
In this work, we focus on a widely-adopted GUI agent paradigm (Lu et al., 2025; Wu et al., 2024c; Qin et al., 2025), specifically a LVLM-powered agent denoted by the model $M$. Initially, the agent receives a system prompt $p_{s}$ and a user instruction $p_{u}$ representing a specific task. At each subsequent interaction step, the agent observes a screenshot $s_{t}$, rendered from the current HTML content, along with the action history $H_{t}$, and then outputs an action $a_{t}$ until the task is completed or failed. Formally, this can be expressed as:
(1) $a_{t}=M(p_{s},p_{u},s_{t},H_{t}),$
where $H_{t}=[a_{1},a_{2},\dots,a_{t-1}]$.
2.3. Environmental Injection Attacks
Environmental Injection Attacks (EIAs) are a critical security concern for GUI agents. These attacks originate externally from manipulated environments, posing a severe threat due to their potential impact. Prior works (Liao et al., 2024; Ma et al., 2024; Lu et al., 2025; Fu et al., 2024; Wang et al., 2025; Yang et al., 2024) have frequently implemented EIAs by directly altering HTML or screenshots. For instance, the EIA (Liao et al., 2024) method injects new form elements into HTML to mislead agents, while ENVINJECTION (Wang et al., 2025) applies adversarial perturbations to entire screenshots. While these methods have demonstrated effectiveness, they typically operate under the assumption that the attacker is a malicious website administrator or developer—an assumption rarely valid in realistic scenarios.
In contrast, we consider a more realistic setting where the attacker is a regular malicious user, such as a content sharer on a social platform, who can only upload a malicious trigger like text or images to the target website. However, many existing methods (Wu et al., 2024b; Zhao et al., 2025) fail to account for dynamic visual contexts, specifically variations in the position and the surrounding content of the trigger. For instance, Wu (Wu et al., 2024b) optimized malicious images without considering any visual background and then deployed them directly onto target websites, severely limiting their attack effectiveness.
Most similar to our work is MIP (Aichberger et al., 2025). While MIP acknowledges dynamic visual context should be considered, it only introduces minor changes to the surrounding content and does not account for variations in the trigger image’s position. Furthermore, the trigger images used in MIP occupy a relatively large area, neglecting the second challenge we highlight in this paper regarding the trigger’s small proportion of total content. In contrast, our work considers significant variations in both trigger positions and surrounding content, closely simulating real-world scenarios characterized by highly dynamic internet content. To our knowledge, this study presents the first systematic investigation into how small trigger images perform within dynamically changing visual contexts when attacking GUI agents.
3. Threat Model
Attack Scenarios. We consider scenarios occurring on online platforms such as social media platforms or e-commerce websites, where users can freely upload images and texts and browse content uploaded by others. Given the large number of users, some may seek to manipulate GUI agents for their personal benefit or to harm other users.
Attacker’s Goal. The attacker aims to optimize a perturbation for a selected trigger image and upload the perturbed image to the target website. When other users browse the platform using GUI agents, the uploaded trigger image induces the agent to execute actions beneficial to the attacker or harmful to other users, such as redirecting users to promotional websites. Specifically, the attacker’s goal is to mislead the agent to produce an incorrect target action when exposed to the trigger image.
Attacker’s Constraints. Attackers face several limitations. They cannot predict the exact position of their uploaded trigger image or the surrounding content when other users access the website. Additionally, attackers have no knowledge of the GUI agent’s action history or the specific user instructions provided to the agent. Moreover, the trigger image typically occupies only a small fraction of the rendered page, further constraining the attack surface.
Attacker’s Capabilities. Attackers can upload trigger images to a target website. Because website layouts and styles typically remain stable over short periods, attackers can reliably anticipate the overall structure and appearance of the target page. Furthermore, we assume that attackers have white-box access to the model powering the GUI agent, including knowledge of its gradients and architecture  [^1] We also consider transfer-based black-box attacks in our evaluation.. Given the increasing deployment of GUI agents and the ease with which trigger images can be distributed online (Wang et al., 2024b; Tur et al., 2025), compromising even one popular model may lead to far-reaching consequences. In practice, even if only a small fraction of users rely on the targeted model, such attacks can still result in significant security risks across the broader internet ecosystem.
4. Our Motivations
(a) Successful case
(b) Unsuccessful case
Figure 1. Attention maps for the two cases. The red box marks the trigger image; warmer colors indicate higher attention. In the successful case, attention is concentrated on the trigger image region, whereas in the unsuccessful case, attention is dispersed across the screenshot.
Under the threat model defined in Section 3, existing environmental injection attacks are either infeasible or exhibit low effectiveness. In this section, we analyze two major challenges faced when attacking GUI agents, motivating our proposed attack method.
Challenge (1) The position and surrounding content of the trigger are unpredictable, making successful attacks in unknown contexts difficult.
Due to the dynamic nature of internet content, attackers cannot anticipate the exact position of their uploaded trigger image or the surrounding visual content. For example, a trigger image could appear at the top-right corner alongside other products of the same category in one user’s view, while in another user’s view, the same trigger image might be centrally placed among best-selling items. Ensuring effective attacks across such diverse and dynamic visual contexts is challenging.
Achieving successful attacks in dynamic environments hinges on acquiring a realistic and diverse set of training data for the target website (Brown et al., 2017). However, manually collecting webpage samples with diverse contexts for training is often laborious and time-consuming (Kasuga and Yonetani, 2024). To overcome this significant hurdle, we introduce LLM-Driven Environment Simulation. This innovative approach leverages the powerful generative capabilities of Large Language Models to automatically synthesize highly realistic and diverse training data. By training the trigger image under these varied contexts, we effectively enhance its generalizability, thereby enabling successful attacks even in dynamic and unknown environments. Detailed descriptions can be found in Section 5.2.
Challenge (2) The trigger image constitutes only a small fraction of the overall website screenshot, making it difficult for the model to consistently focus on it.
Although our experiments show that the attack effectiveness improves substantially when LLM-Driven Environment Simulation is employed during training, the overall performance still falls short of expectations. We attribute this to the limited size of the trigger image relative to the entire webpage screenshot, typically constituting around 5%, thereby limiting the trigger image’s ability to consistently affect the agent’s decision-making process.
We hypothesize that consistently directing the model’s attention to the trigger region is crucial for overcoming this challenge (Lu et al., 2025). Empirical results obtained from evaluating trigger images trained solely with LLM-Driven Environment Simulation support this hypothesis. Specifically, in about 82% of successful attack cases, the average attention weight on the trigger image region is higher than that in other regions. In contrast, this proportion drops to roughly 61% in failed cases, a value that is close to a random distribution. Figure 1 also exemplifies this, showing significant attention concentration on the trigger region in successful attack case, while attention is dispersed evenly across the screenshot in unsuccessful case.
Motivated by these insights, we introduce Attention Black Hole, which explicitly encourages the model’s attention to converge on the trigger image region. This technique significantly enhances attack effectiveness, effectively addressing the second challenge. Further details can be found in Section 5.3.
5. Methodology
To address the above challenges, we propose Chameleon, an effective approach for attacking existing GUI agents. We first detail the overall attack procedure. Then, we provide an in-depth explanation of our LLM-Driven Environment Simulation, which enables the automatic generation of realistic and dynamic training data. Finally, we introduce Attention Black Hole, designed to encourage the model’s attention to converge on the trigger image region. The overview of Chameleon is depicted in Figure 2.
Figure 2. Overview of our proposed Chameleon.
5.1. Overview
For a selected target website, such as a social media platform like RedNote  [^2] https://www.xiaohongshu.com/ or an e-commerce site like Amazon  [^3] https://www.amazon.com/, an attacker first chooses a raw trigger image $I$, such as a product image or a post cover that appears normal and benign, which will then be optimized during subsequent training.
At each training step, we first employ our proposed LLM-Driven Environment Simulation $G_{LES}$ (Section 5.2), to sample a new realistic context. This provides a webpage screenshot $s$ embedded with the perturbed trigger image, a corresponding user instruction $p_{u}$, and a pixel-level mask $mask^{shot}$ that identifies the trigger’s location within $s$. The perturbation $\delta$ within the trigger image region is then updated using Projected Gradient Descent (PGD) (Madry et al., 2017). To ensure the resulting malicious trigger image remains imperceptible to users, we constrain the perturbation $\delta$ to a predefined $\ell_{\infty}$ norm bound $\epsilon$, formally:
(2) $|\delta|_{\infty}\leq\epsilon.$
[…]
(3) $\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{CE}}(a,\hat{a})+\lambda\cdot\mathcal{L}_{\mathrm{attn}},$
where $\lambda$ is a hyperparameter used to balance the influence between these two loss terms.
After the training procedure described above, the attacker uploads the resulting trigger image to the target website, acting as a regular user. Once the trigger image is integrated into a public post or product listing, any GUI agent browsing the site and encountering the trigger image within a webpage will be susceptible to attack, potentially performing the intended harmful action. Algorithm 1 summarizes the overall procedure of Chameleon, covering the preprocessing, training, and deployment stages.
Input: Target website URL, raw trigger image $I$, GUI agent $M$, system prompt $p_{s}$, target action $\hat{a}$, perturbation bound $\epsilon$, step size $\alpha$, weight $\lambda$, training steps $K$, flagship LLM
/* Preprocessing */
$T\leftarrow$ Parse target website with flagship LLM;
  /* HTML template */
$C\leftarrow$ Crawl large-scale multi-modal content;
  /* image-text corpus */
/* Training */
1 Initialize perturbation $\delta\leftarrow 0$ ;
2 for $k=1$ to $K$ do
  /* LLM-Driven Environment Simulation (LES) */
3  $(s,mask,p_{u})\leftarrow G_{LES}(T,C,I+\delta)$ ;
4  $(a,A)\leftarrow M(p_{s},p_{u},s)$ ;
  /* Attention Black Hole (ABH) */
5  $\mathcal{L}_{attn}\leftarrow\textsc{ABH}(A,mask)$ ;
[…]
7  $\mathcal{L}\leftarrow\mathcal{L}_{CE}+\lambda\mathcal{L}_{attn}$ ;
8  $\delta\leftarrow\text{Update}(\delta,\mathcal{L},\epsilon,\alpha)$ ;
9
10 end for
/* Deployment */
11 $I_{adv}\leftarrow I+\delta$ ;
12 Upload $I_{adv}$ to target website ;
Algorithm 1 Chameleon
5.2. LLM-Driven Environment Simulation
To address the challenge posed by the variability in trigger positions and their surrounding content, we introduce LLM-Driven Environment Simulation (LES). This approach leverages known webpage layouts and styles, employing large language models to automatically synthesize realistic and diverse training samples that capture authentic contexts. Specifically, we define a generation function $G_{LES}$, which outputs a rendered webpage screenshot, a binary mask indicating the location of the trigger image, and contextually relevant user instructions for downstream training.
First, to ensure our simulated environments are realistic, we create a high-fidelity HTML template $T$ for each target website. We leverage a flagship LLM to parse the live website, preserving its core structural and stylistic components while removing existing content. This results in a template that is visually authentic.
We then construct a large-scale, multi-modal corpus $C$ by crawling the target website for actual content, gathering an average of over 5,000 distinct items (e.g., product images, titles, prices) for each website. This process ensures our content pool is not only realistic but also sufficiently diverse to simulate a wide range of scenarios.
Next, we define an HTML generation function $g$, which integrates randomly selected contextual image-text pairs from corpus $C$ along with the pre-selected trigger image $I$ into the HTML template $T$, producing renderable HTML code $h$:
(4) $h=g(T,C,I).$
The generated HTML code is subsequently rendered into a screenshot $s$ accompanied by a binary pixel-level mask $mask^{shot}$ marking the trigger region, where pixels corresponding to the trigger image region are set to 1, and all remaining pixels are set to 0:
(5) $s,mask^{shot}=f(h).$
Considering the inherently unpredictable nature of real user instructions during actual usage, we utilize another advanced LLM to automatically generate realistic and diverse user instructions $p_{u}$ corresponding to each synthesized screenshot $s$. Thus, the complete environment generation process within our LES framework can be formally described as:
(6) $s,mask^{shot},p_{u}=G_{LES}(f(g(T,C,I))).$
Through this comprehensive procedure, LES effectively combines known webpage structures with realistic and diverse environment generation, thereby significantly enhancing the generalizability of optimized trigger images under dynamically varying web environments.
5.3. Attention Black Hole
To address the challenge arising from the small proportion of the trigger image within the overall webpage screenshot, we produce Attention Black Hole (ABH), which optimizes the trigger image to effectively attract the model’s attention toward the trigger region.
Suppose that after passing through the visual encoder and connector, the input screenshot is transformed into a sequence of image tokens with length $n\times m$. Given the text tokens comprising system prompt $p_{s}$, user instruction $p_{u}$, and action history $H_{t}$ along with image tokens, the LVLM generates a new token sequence $N$. We utilize attention weights from the last layer of the LVLM to quantify the interactions between each newly generated token and all image tokens, producing an attention map $A_{i,j}$ defined as follows:
(7) $A_{i,j}=\frac{1}{|N|\times H}\sum_{t=1}^{|N|}\sum_{h=1}^{H}Attention_{t,h}^{i,j},$
where $Attention_{t,h}^{i,j}$ denotes the attention weight from the $h$-th attention head, relating the $t$-th new token to the $(j+m\times(i-1))$-th image token. $H$ represents the number of attention heads, and $|N|$ indicates the total number of newly generated tokens.
Based on the binary mask $mask^{shot}$ of the trigger image at the pixel level, we apply a resizing operation to obtain a token-level binary mask $mask^{attn}$ of size $n\times m$, where $mask^{attn}_{i,j}$ is set to 1 if the $(j+m\times(i-1))$-th image token corresponds to a patch that overlaps with the trigger image region, and 0 otherwise.
Subsequently, we define $\overline{A}_{in}$, representing the model’s degree of focus on the trigger image region, as the mean attention weight within the trigger image region. Likewise, $\overline{A}_{out}$ represents the model’s attention to other regions of the screenshot and is computed as the mean attention weight outside the trigger region. Formally, these are given by:
(8) $\overline{A}_{in}=\frac{\sum_{i,j}A_{i,j}\times mask^{attn}_{i,j}}{\sum_{i,j}mask^{attn}_{i,j}},$
(9) $\overline{A}_{out}=\frac{\sum_{i,j}A_{i,j}\times(1-mask^{attn}_{i,j})}{\sum_{i,j}(1-mask^{attn}_{i,j})}.$
Finally, we define the loss function $\mathcal{L}_{attn}$ as the ratio of the average attention outside of the trigger region to the average attention within it. Formally, $\mathcal{L}_{attn}$ is defined as:
(10) $\mathcal{L}_{attn}=\frac{\overline{A}_{out}}{\overline{A}_{in}}.$
Minimizing this loss explicitly guides the model to focus on the trigger region, which we regard as the key to overcoming the challenge posed by the small size of the trigger image.
6. Experiments
To systematically evaluate our proposed Chameleon, we conduct extensive experiments designed to answer the following Research Questions (RQs).
First, as stated in the threat model, the primary goal of an attacker is to mislead the GUI agent into executing an incorrect action when exposed to the trigger image. We therefore design our first research question to evaluate the fundamental effectiveness of Chameleon in achieving this objective.
RQ1: How effective is Chameleon for conducting environmental injection attacks against GUI agents? To answer this, we evaluate the attack success rates of Chameleon on six distinct and well-known target websites.
Although compromising even a single popular model can pose a significant security risk due to the vast user base of the internet, an attacker’s impact further increases if a trigger crafted against one LVLM transfers to unseen LVLMs. We therefore focus exclusively on cross-model transferability.
RQ2: How transferable is Chameleon to unseen LVLMs? We train triggers on a surrogate LVLM and evaluate them—without any adaptation—on multiple target LVLMs to measure cross-model transferability.
Table 1. Datasets used in this work. Trigger Coverage Ratio denotes the percentage of the screenshot covered by the trigger image for each website.
Category Website URL Trigger Coverage Ratio
Shopping Amazon https://www.amazon.com/ 4.67%
Taobao https://www.taobao.com/ 7.59%
Social Media RedNote https://www.xiaohongshu.com/ 8.19%
Bilibili https://www.bilibili.com/ 9.97%
Music Streaming NetEase Cloud Music http://ir.music.163.com/en/ 8.17%
QQ Music https://y.qq.com/ 5.99%
Finally, we aim to understand the individual contributions of our core technical innovations. We therefore design our third research question to analyze the importance of each component within our framework.
RQ3: How do the LLM-Driven Environment Simulation and Attention Black Hole contribute to the performance of Chameleon? We perform an ablation study to isolate and quantify the impact of each of these components on the overall effectiveness of the attack.
6.1. Experimental Setup
Datasets. To ensure the realism of our evaluation, we selected six target websites spanning three representative categories of GUI-based online services, as summarized in Table 1. These categories reflect diverse user interaction patterns and GUI design paradigms, which are critical for assessing the robustness of Chameleon across practical use cases. For each website, we construct a validation set of 300 and a test set of 1,200 screenshot-instruction pairs using the LLM-Driven environment simulation introduced in Section 5.2. Importantly, the images and instructions in the training, validation, and test sets are mutually disjoint, ensuring rigorous evaluation. Figure 3 displays one representative screenshot for each of the six websites.
Figure 3. Representative screenshots for each website. Trigger images are outlined in red.
LVLMs for GUI agents. Our evaluation is conducted on a diverse set of four popular Large Vision-Language Models. These include two models specialized for GUI tasks, UI-TARS-7B-DPO (Qin et al., 2025) and OS-Atlas-Base-7B (Wu et al., 2024c), as well as two general-purpose LVLMs, Qwen2-VL-7B (Wang et al., 2024a) and LLaVA-1.5-13B (Liu et al., 2023). Notably, UI-TARS-7B-DPO and OS-Atlas-Base-7B are fine-tuned from Qwen2-VL-7B.
Baselines. We adopt a standard PGD-based adversarial attack as our baseline (Madry et al., 2017), representing methods utilized in previous research (Wu et al., 2024b; Aichberger et al., 2025). During the optimization of the trigger image, this baseline approach does not incorporate any environmental context, such as surrounding webpage content or stylistic information.
Target Action. Consistent with the attacker’s goal, we define the target action as instructing the agent to navigate to a specific malicious URL. This could be a URL for a promotional site or a more harmful phishing page. For each target website, we set the target action to a navigation command that directs the agent to a malicious URL such as:
goto [http://malicious-example.com]
Evaluation Metrics. We employ the Attack Success Rate (ASR) to evaluate the effectiveness of Chameleon, formally defined as follows:
(11) $\text{ASR}=\frac{N_{\text{attack}}}{N_{\text{total}}},$
where $N_{\text{attack}}$ denotes the number of responses matching the target action within the test set, and $N_{\text{total}}$ represents the total size of the test set. A strict string matching criterion is adopted to verify whether the agent’s response matches the target action precisely. Consequently, our ASR metric provides a conservative estimate, representing a lower bound for the capability of the trigger image to pose practical security threats.
Implementation Details. We employ GPT-4o (Achiam et al., 2023) to parse the live website, preserving its core structural and stylistic components while removing existing content, and we use Qwen2.5-VL-32B-Instruct (Bai et al., 2025) to automatically generate realistic and diverse user instructions. The system prompt and user prompt for the GUI agent, as well as the agent’s action space, are adapted from VisualWebArena (Koh et al., 2024). Since it is challenging to collect authentic action histories, we follow ENVINJECTION (Wang et al., 2025) and randomly sample 0 to 10 historical actions for each instance. To ensure rigorous evaluation, we maintain a strict separation, with no overlap in the sampled action histories used across the training, validation, and testing. The hyperparameter $\lambda$ is set to $0.3$ and $\epsilon$ is set to $\frac{32}{255}$. The perturbation is optimized for a total of $5,000$ steps, with updates performed at each step using a fixed step size of $\alpha=\frac{1}{255}$. All experiments are conducted on a server equipped with 8 NVIDIA A100-PCIE-40GB GPUs.
6.2. RQ1: Effectiveness in Attacking Target GUI Agents
In this RQ, we evaluate whether Chameleon effectively misleads GUI agents into executing a predefined malicious action when exposed to an adversarial trigger image embedded within a target website.
Setup. We compare the ASR of Chameleon against the PGD baseline across four GUI agents on the six target websites detailed in our experimental setup.
Table 2. ASRs (%) of baseline and Chameleon across different GUI agents and websites. NetEase is used as an abbreviation for NetEase Cloud Music. The percentage shown at the top-right of each website name denotes the trigger coverage ratio (percentage of screenshot pixels occupied by the trigger image).
Model Method Shopping Social Media Music Streaming Avg.
Amazon4.67% Taobao7.59% RedNote8.19% Bilibili9.97% NetEase8.17% QQ Music5.99%
UI-TARS-7B-DPO Baseline 3.17 2.83 4.08 5.75 6.58 1.25 3.94
Chameleon 20.75 22.58 23.25 41.58 41.08 8.08 26.22
OS-Atlas-Base-7B Baseline 4.25 5.17 3.83 7.08 8.00 3.25 5.26
Chameleon 23.67 33.17 26.42 35.50 43.75 33.08 32.60
Qwen2-VL-7B Baseline 2.50 4.00 4.67 6.83 7.58 2.08 4.61
Chameleon 9.42 14.42 17.58 22.42 25.93 9.92 16.62
LLaVA-1.5-13B Baseline 4.75 6.25 7.08 8.17 8.75 3.50 6.42
Chameleon 37.17 43.00 60.75 50.83 76.58 32.08 50.07
Results. Table 2 presents the detailed ASR results for both our method and the baseline across all experimental configurations. Higher ASR indicates stronger attack effectiveness.
Analyses.Chameleon consistently outperforms the baseline by a substantial margin across all experimental configurations. While prior work (Wu et al., 2024b) has reported relatively high attack success rates for baseline methods when applied to static websites, Table 2 shows that the same baseline performs poorly in dynamically varying visual contexts, with ASRs remaining close to zero. In contrast, Chameleon achieves considerably higher ASRs across all settings, demonstrating its strong ability to reliably induce malicious behaviors in GUI agents under dynamic conditions. ❷ The effectiveness of Chameleon is positively correlated with the relative size of the trigger image within the screenshot. On websites such as NetEase Cloud Music, where the trigger image constitutes a larger portion of the visual area, the effectiveness is substantially higher, often exceeding 50%. Conversely, on websites such as Taobao, where the trigger image occupies a smaller region, the effectiveness remains below 20%. This finding is consistent with our earlier observation that when the trigger image represents only a small fraction of the overall screenshot, it becomes more difficult for the model to consistently attend to it.
Answer to RQ1: Chameleon demonstrates strong effectiveness in misleading GUI agents and consistently achieves substantially higher ASRs than baseline methods.
6.3. RQ2: Generalization Ability Across LVLMs
In this RQ, we evaluate the generalization ability of our trigger image across multiple LVLMs. Specifically, we aim to examine whether a trigger trained on one model can generalize to other unseen models in a black-box setting.
Setup. We conduct a $4\times 4$ transfer experiment across four LVLMs: UI-TARS-7B-DPO, OS-Atlas-Base-7B, Qwen2-VL-7B and LLaVA-1.5-13B. For each surrogate model, we optimize the trigger with white-box access to that model only, then evaluate it zero-shot on the remaining three target models under black-box access. Transferability is quantified by the ASR.
(a) NetEase Cloud Music
(b) Amazon
(c) RedNote
Figure 4. Transferability of Chameleon across models. Each cell represents the ASR (%) when the trigger image is trained on the surrogate model (row) and tested on the target model (column).
Results. The cross-model transferability results are shown in Figure 4. Due to space limitations, we present results only on three representative target websites, namely NetEase Cloud Music, Amazon, and RedNote.
Analyses.Transfer is stronger between related models. OS-Atlas-Base-7B and UI-TARS-7B-DPO are both fine-tuned from Qwen2-VL-7B, and we observe moderate bidirectional transfer within this family. For instance, Qwen2-VL-7B $\to$ OS-Atlas-Base-7B achieves 14.11% and OS-Atlas-Base-7B $\to$ Qwen2-VL-7B achieves 19.17%. These results indicate that shared architectures and training data lead to overlapping representations, which can be exploited by adversarial triggers. This observation underscores a practical security concern: a trigger optimized against one model can remain effective against other closely related variants, thereby broadening the potential impact of our proposed Chameleon. ❷ Transfer collapses across dissimilar models. In contrast, LLaVA-1.5-13B differs substantially from the Qwen2-VL family in both architecture and training data. As a result, all transfer pairs involving LLaVA-1.5-13B yield 0.00% ASR in our experiments. This sharp contrast highlights the limits of transferability when model families differ substantially, suggesting that architectural heterogeneity can serve as a natural barrier against cross-model attacks, although it does not eliminate risks in the widely used internet ecosystem.
Answer to RQ2: Chameleon transfers well between similar models but shows negligible transfer across dissimilar ones (e.g., involving LLaVA-1.5-13B).
6.4. RQ3: Ablation Study
In this RQ, we investigate the contribution of the LLM-Driven Environment Simulation (LES) and the Attention Black Hole (ABH) to the effectiveness of Chameleon.
Setup. We conduct an ablation study on OS-Atlas-Base-7B and LLaVA-1.5-13B by evaluating three settings across six websites: ❶ Chameleon : the full approach with both LES and ABH; ❷ Removing LES: the trigger image is trained on a limited set of 100 manually collected screenshots, without the automatic and scalable context construction provided by LES; ❸ Removing ABH: the trigger image is trained with only the cross-entropy loss $\mathcal{L}_{\mathrm{CE}}$, omitting $\mathcal{L}_{\mathrm{attn}}$.
Table 3. ASRs (%) under different ablation settings. NetEase is used as an abbreviation for NetEase Cloud Music. The percentage shown at the top-right of each website name denotes the trigger coverage ratio (percentage of screenshot pixels occupied by the trigger image).
Method Shopping Social Media Music Streaming Avg.
Amazon4.67% Taobao7.59% RedNote8.19% Bilibili9.97% NetEase8.17% QQ Music5.99%
OS-Atlas-Base-7B
Chameleon 23.67 33.17 26.42 35.50 43.75 33.08 32.60
w/o LES 14.89 (-8.78) 15.26 (-17.91) 13.28 (-13.14) 20.81 (-14.69) 20.11 (-23.64) 15.73 (-17.35) 16.68 (-15.92)
w/o ABH 17.97 (-5.7) 29.59 (-3.58) 24.10 (-2.32) 30.02 (-5.48) 39.77 (-3.98) 26.76 (-6.32) 28.20 (-4.4)
LLaVA-1.5-13B
Chameleon 37.17 43.00 60.75 50.83 76.58 32.08 50.07
w/o LES 20.75 (-16.42) 29.75 (-13.25) 34.42 (-26.33) 35.67 (-15.16) 60.50 (-16.08) 23.09 (-8.99) 34.03 (-16.04)
w/o ABH 26.83 (-10.34) 34.92 (-8.08) 53.50 (-7.25) 46.71 (-4.12) 73.00 (-3.58) 20.75 (-11.33) 42.62 (-7.45)
Results. The detailed results for ablation study are presented in Table 3.
Analyses.Contribution of LES. Removing LES leads to a consistent reduction in ASR across all websites. For example, with OS-Atlas-Base-7B on NetEase Cloud Music, the ASR decreases from 43.75% to 20.11% when LES is removed. Similarly, with LLaVA-1.5-13B on RedNote, the ASR drops from 60.75% to 34.42%. These results indicate that large-scale automatic simulation of realistic and diverse visual context during training is critical to achieving effective attacks. ❷ Contribution of ABH. Removing ABH also reduces ASR on every website. Notably, the largest declines appear when the trigger image occupies a smaller portion of the screenshot. For instance, with OS-Atlas-Base-7B on QQ Music, where the trigger covers 5.99% of the screenshot, the ASR decreases by 6.32%, while on RedNote, with a coverage of 8.19%, the decrease is only 2.32%. For LLaVA-1.5-13B, the ASR drops by 10.34% on Amazon, where the trigger coverage is 4.67%, but only 4.12% on Bilibili, where the coverage reaches 9.97%. We hypothesize that this is because, when the trigger image occupies a larger portion of the screenshot, the model can more readily focus on the trigger image region even without explicit guidance.
Answer to RQ3: Both LES and ABH are necessary. LES improves robustness by providing automatic simulation of realistic and diverse contexts, while ABH amplifies the trigger’s effect by explicitly directing attention to the trigger region.
7. Case Study
Figure 5. Case study: GUI agent’s behavior for the task “open the first playlist in the homepage recommendations of NetEase Cloud Music.” (a) Homepage with the uploaded trigger image. (b) Without attack: correct navigation to the first playlist. (c)With attack: incorrect navigation to a promotional site.
To assess whether Chameleon can induce harmful actions in a realistic interactive setting, we conducted a closed-loop evaluation that more closely reflects real-world scenarios.
Experimental Setup. We instantiated a GUI agent using UI-TARS-7B-DPO as the underlying LVLM. The agent ran end-to-end without human assistance, issuing actions that directly controlled a browser. To avoid any risk to production services, all experiments were performed in a fully isolated sandbox. Specifically, we constructed a static website that mirrors the visual layout and interaction patterns of NetEase Cloud Music, deployed on localhost inside a containerized network where all outbound traffic was blocked at the firewall and DNS resolution was disabled. The agent’s initial task was: “open the first playlist in the homepage recommendations of NetEase Cloud Music.” The agent launched the browser, entered the public NetEase Cloud Music URL as a string, and the request was internally rewritten to the local mirror, which returned the sandboxed homepage. We instrumented the environment to record screenshots and the agent’s actions at every step.
The homepage contained a single trigger image embedded in the recommendation rail. We evaluated two conditions that were identical in layout and ordering of items. In the without attack condition, the trigger image was the original clean image. In the with attack condition, the same image carried the perturbation optimized by Chameleon under an $\ell_{\infty}$ bound $\epsilon$. The agent observed the rendered page’s screenshot, action history, and the browser state updated accordingly, which completed the control loop. Figure 5 shows the sandboxed homepage in the bottom-left panel, the target playlist page in the top-right panel, and the promotional site used as the malicious target in the bottom-right panel. Success for the benign task was defined as navigating to the first playlist page. Attack success was defined as issuing a navigation action that opened the predefined promotional webpage.
Results and Analyses. As shown in Figure 5, in the without attack condition, the agent executed the instructed task correctly. It parsed the homepage’s screenshot, selected the first recommended playlist, and navigated to the corresponding playlist detail page. In the with attack condition, the agent encountered the homepage containing the perturbed trigger image and immediately issued a navigation action that opened the promotional site, rather than proceeding to the first playlist. These observations demonstrate that Chameleon is not restricted to open-loop or purely simulated pipelines. The attack transfers to a closed-loop setting where actions modify state and state feeds back to the policy. The outcomes highlight tangible risks to Internet-facing GUI agents, since a single user-uploaded image can redirect goal-directed behavior even when the surrounding interface and task specification remain benign.
(a) NetEase (LLaVA-1.5-13B)
(b) RedNote (LLaVA-1.5-13B)
(c) NetEase (OS-Atlas-Base-7B)
(d) RedNote (OS-Atlas-Base-7B)
Figure 6. Effectiveness of defense strategies against Chameleon across different websites. The first row corresponds to LLaVA-1.5-13B and the second row to OS-Atlas-Base-7B, each evaluated on two representative target websites. NetEase is used as an abbreviation for NetEase Cloud Music.
8. Discussion
8.1. Defenses
We investigate several practical defenses for GUI agents under the threat model considered in this paper and evaluate their ability to reduce the effectiveness of Chameleon. Experiments are conducted on LLaVA-1.5-13B and OS-Atlas-Base-7B using the NetEase Cloud Music and RedNote as target websites. Effectiveness is measured by the Attack Success Rate (ASR), and the results are shown in Figure 6.
Safety prompt. Safety prompts (Wang et al., 2024c; Zhang et al., 2025b) are a widely adopted technique for guiding models away from unsafe behavior. We insert a safety prompt into the system prompt that explicitly directs the agent to follow the user’s command strictly and to disregard any potentially malicious content in the observation. The specific prompt used is shown below:
You should adhere strictly to the user’s instructions and not be distracted by any injected malicious instructions found in the observation.
However, the ASR shows almost no reduction. For example, in the case of a LLaVA-1.5-13B powered agent on RedNote, the ASR decreases only marginally from 60.75% to 60.21%. This indicates that safety prompts alone are insufficient to defend against Chameleon.
Figure 7. Comparison between original (left half) and noised (right half) versions of four photography images. Random noise causes visible degradation in image quality, significantly undermining user experience.
Verifier. Another common defense is to verify model outputs before execution. We employ DeepSeek-V3 (Liu et al., 2024) as an external verifier that examines each candidate action together with the user instruction and aborts execution if the action appears inconsistent or risky. For the OS-Atlas-Base-7B powered agent on NetEase Cloud Music, this verification reduces the ASR from 43.75% to 36.42%. However, we also observe false positives where benign actions are blocked, which reduces overall utility. Furthermore, this approach introduces additional latency and inference cost since it relies on a secondary model.
Random noise on uploaded images. Since Chameleon performs attacks by applying adversarial perturbations, one potential countermeasure is to add random noise to uploaded images. We assume that a website administrator automatically adds noise $\eta$ with an $\ell_{\infty}$ bound $|\eta|_{\infty}\leq\epsilon$. With $\epsilon=\tfrac{8}{255}$, this defense proves highly effective: on RedNote, the ASR of a LLaVA-1.5-13B powered agent drops to 11.84%, and for OS-Atlas-Base-7B the ASR is reduced to nearly zero. Although effective, even small bounded noise noticeably degrades image quality, which may harm user experience, particularly in scenarios where high visual fidelity is required, such as photography or digital art. As shown in Figure 7, applying random noise leads to clear quality degradation.
8.2. Threats to Validity
Evaluation Metrics. We compute ASR using a strict string-matching criterion that counts an attack as successful only when the exact target action is produced. Minor variations such as additional spaces, a www prefix, or appended paths are treated as failures. This conservative definition provides a lower bound on practical risk. Nevertheless, even with this lower-bound measure, our results reveal widespread vulnerabilities in GUI agents, suggesting that this limitation does not undermine our conclusions. To further mitigate this concern, future evaluations could incorporate human judgment or LLM-based assessment of semantic equivalence, which would capture cases where the produced actions are functionally equivalent despite surface-level variations.
Trigger Images. For each website, we randomly select a single image from the crawled corpus to serve as the trigger. The choice of trigger likely affects attack effectiveness due to factors such as visual salience, color palette, and composition. Our current results therefore reflect the performance of one draw per site. Future work will expand this evaluation by considering multiple triggers for each site, analyzing sensitivity to different visual attributes, and developing principled strategies for selecting images that maximize attack effectiveness. A more systematic study of trigger characteristics would also provide insights into which types of content are more likely to compromise GUI agents in practice.
Datasets. Existing GUI agent datasets often present webpages where text and images are static or where images occupy a disproportionately large fraction of the webpage, which misaligns with real-world internet scenarios. To address this, we construct test sets using our LLM-Driven Environment Simulation, which enables large-scale generation of realistic and diverse visual contexts. Moreover, to eliminate potential data leakage, we ensure that the screenshots used in the training, validation, and test sets do not share any uploaded images. This separation guarantees that performance improvements cannot be attributed to memorization of specific samples.
Replication of Our Experiments. The behavior of LVLM-powered agents can be influenced by multiple factors, such as decoding temperature, which complicates replication. To support reproducibility, we release detailed descriptions of our experimental settings, including hyperparameters and environment configurations, along with links to the exact model checkpoints used. In addition, our full code repository is publicly available to facilitate independent verification. These measures collectively enhance transparency and provide a reliable foundation for reproducing our findings. Future work may also consider standardized benchmarks and controlled evaluation environments, which would further reduce randomness and strengthen replicability.
Ethics Statement. In this work, we propose a more practical and realistic threat model and an effective attack approach. However, our goal is not to promote malicious behavior, but rather to reveal the vulnerabilities of widely used GUI agents when deployed in real-world, open-ended internet environments. By exposing these vulnerabilities, we aim to raise awareness of the potential risks and emphasize the urgent need for robust and practical defenses. We hope that our findings will inform future research on building safer and more trustworthy web-based agent systems.
9. Conclusion
In this paper, we present a more realistic threat model in which the attacker is a regular user who can only upload a small trigger image that appears within a dynamically changing environment. To address the challenges posed by dynamic visual contexts and limited screenshot coverage, we propose Chameleon , a novel attack framework that introduces two key novelties:LLM-Driven Environment Simulation, which enables large-scale automatic generation of realistic and diverse webpage simulations, and Attention Black Hole, which explicitly guides the agent’s focus toward the trigger region. Extensive experiments across multiple websites and models demonstrate that Chameleon significantly outperforms existing methods in attack success rate. Ablation and closed-loop evaluations further confirm the effectiveness and real-world applicability of the proposed techniques. Through an evaluation of several commonly used defense strategies, we find that existing defense strategies fail to effectively mitigate the threat of Chameleon without compromising the utility of the agent. Overall, our study reveals the inherent vulnerabilities in widely used LVLM-powered GUI agents. Future work may explore automated trigger detection and effective defenses that preserve usability in open-world web environments.
10. Data Availability
The source code and access details for the datasets can be accessed through an anonymous repository at: https://github.com/zhangyitonggg/attack2gui.