### v2

# Environmental Injection Attacks against GUI Agents in Realistic Dynamic Environments

Abstract.
Graphical User Interface (GUI) agents are increasingly deployed to interact with online web services, yet their exposure to open-world content renders them vulnerable to Environmental Injection Attacks (EIAs). In these attacks, an attacker can inject crafted triggers into website to manipulate the behavior of GUI agents used by other users. In this paper, we find that most existing EIA studies fall short of realism. In particular, they fail to capture the dynamic nature of real-world web content, often assuming that a trigger’s on-screen position and surrounding visual context remain largely consistent between training and testing. To better reflect practice, we introduce a realistic dynamic-environment threat model in which the attacker is a regular user and the trigger is embedded within a dynamically changing environment. Under this threat model, existing approaches largely fail, suggesting that their effectiveness in exposing GUI agent vulnerabilities has been substantially overestimated.
To expose the hidden vulnerabilities of existing GUI agents effectively, we propose Chameleon, an attack framework with two key novelties designed for dynamic environments. ❶ To synthesize more realistic training data, we introduce LLM-Driven Environment Simulation, which automatically generates diverse, high-fidelity webpage simulations that mimic the variability of real-world dynamic environments. ❷ To optimize the trigger more effectively, we introduce Attention Black Hole, which converts attention weights into explicit supervisory signals. This mechanism encourages the agent to remain insensitive to irrelevant surrounding content, thereby improving robustness in dynamic environments. We evaluate Chameleon on six realistic websites and four representative LVLM-powered GUI agents, where it significantly outperforms existing methods. Ablation studies confirm that both novelties are critical to performance, and a closed-loop sandbox experiment further demonstrates that Chameleon can successfully hijack agent behavior in conditions that closely mirror real-world usage. Our results uncover a critical, previously underexplored vulnerability of GUI agents in realistic dynamic environments and establish a robust foundation for future research on defense in open-world GUI agent systems.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
With the rapid development of Large Language Models (LLMs) and Large Vision-Language Models (LVLMs) (Achiam et al., 2023; Wang et al., 2024a; Bai et al., 2025; Wu et al., 2024c; Nong et al., 2024), a new class of powerful agents called GUI agents has been designed to interact with Graphical User Interfaces (GUIs) (Hong et al., 2024; Qin et al., 2025; Wu et al., 2024b; Zhang et al., 2024a). These GUI agents can execute increasingly sophisticated tasks, moving beyond simple dialogue to perform complex operations like website manipulation (Zhou et al., 2023). A key characteristic of these agents is their ability to autonomously access and act upon live internet content. While this capability greatly expands their functionality, it simultaneously introduces novel security risks inherent to open-world interaction (Nguyen et al., 2024; Wang et al., 2024b; Johnson et al., 2025; Yang et al., 2025a, b; Ye et al., 2025b; Chen et al., 2025).
A growing body of work have identified that GUI agents are particularly vulnerable to Environmental Injection Attacks (EIAs), which originate from the external environment rather than from the user themselves (Liao et al., 2024; Ma et al., 2024; Aichberger et al., 2025; Zhang et al., 2024c). Recently, EIA research has increasingly adopted a practical assumption that the attacker is a malicious regular user without administrative privileges, and can only influence the webpage through normal content uploads. For instance, on an e-commerce platform, a malicious attacker could upload a specially crafted trigger image disguised as a product photo. When another user’s GUI agent encounters this trigger while browsing the website, it might be automatically induced to navigate to a specific promotional site, all without the user’s explicit instruction (Zhao et al., 2025; Lu et al., 2025).
However, we find that most existing EIAs still fall short of realism. In particular, prior work (Wu et al., 2024a; Aichberger et al., 2025; Zhao et al., 2025) often evaluates triggers under static or only slightly varying environments, implicitly assuming that the trigger’s on-screen position and surrounding visual context remain largely consistent between training and testing. This assumption conflicts with how modern websites operate. In practice, layouts and nearby content change continuously due to dynamic ranking and recommendation updates, advertisement placement, and frequent content refresh. To illustrate this gap, consider a common e-commerce scenario. Figure 1 shows two screenshots from JD.com  [^1] https://search.jd.com/, collected from two separate sessions that issue the same query (“Apple”). Even in this simple setting, the trigger image highlighted in red appears at different positions on the page, and its surrounding context, including nearby products, banners, and text snippets, also changes substantially. Such environmental dynamics pose a fundamental challenge to existing attacks. Many EIA methods optimize and evaluate trigger images in static or slightly varying environments. When deployed in realistic dynamic environments, shifts in trigger position and changes in surrounding visual context may sharply reduce attack effectiveness. These observations motivate a novel threat model that explicitly accounts for environmental dynamism, rather than relying on overly idealistic static assumptions.
Figure 1. An illustration of the dynamic environment in realistic GUI agent application scenarios. The two screenshots show consecutive searches for “Apple” on the same e-commerce platform. We assume the image highlighted by the red box is a trigger uploaded by a attacker.
In this paper, we formally define a realistic dynamic-environment threat model, in which the attacker can hijack GUI Agent by upload trigger images as a regular user. Crucially, these triggers are embedded within a dynamically changing environment, where their positions and surrounding visual context may continuously change and are largely beyond the attacker’s control. Under this threat model, our experiments in Section 5.2 show that trigger images optimized using existing environmental injection attack approaches exhibit near-zero attack success rates (Aichberger et al., 2025; Wu et al., 2024a; Zhao et al., 2025; Madry et al., 2017). This is largely because prior methods optimize triggers under static or slightly varying environments, and thus fail to generalize to dynamic environments. Such a mismatch can mislead security analysis and hinder the development of effective defenses for GUI agents in realistic deployments.
To more fully expose the vulnerabilities of existing GUI agents in realistic deployments, we propose Chameleon. This novel attack framework is specifically designed to ensure that trigger images remain effective within dynamic environments, where on-screen positions and surrounding visual contexts vary unpredictably. Chameleon addresses two key challenges. ❶ How to synthesize large-scale training data that captures dynamic environments? Achieving robust attacks in dynamic environments requires a large number of realistic training samples that cover diverse trigger placements and contextual variations. However, manually collecting webpage samples with diverse layouts and surrounding content is laborious and time-consuming (Koh et al., 2024; Liu et al., 2025; Kasuga and Yonetani, 2024). To overcome this challenge, we introduce LLM-Driven Environment Simulation. Leveraging the generative capabilities of LLMs, we automatically construct high-fidelity simulations of target websites and systematically vary both the trigger placement and its surrounding context. This process produces a large volume of realistic screenshots, enabling the optimized trigger to generalize effectively across dynamic environments. ❷ How to train the trigger more effectively? We find that even with sufficient realistic training data, attack performance can remain suboptimal. Based on our preliminary observations, the agent’s attention is often distracted by constantly changing environments, which can lead to suboptimal performance even with sufficient training data. As shown in Figure 2, successful attacks occur when attention remains concentrated on the trigger, whereas unsuccessful cases arise when attention is dispersed across other salient interface elements. To address this issue, we propose Attention Black Hole, which enables more effective trigger optimization in dynamic environments. The key idea is to explicitly guide the model to focus on the trigger region while suppressing interference from non-trigger areas. Concretely, Attention Black Hole introduces an explicit supervisory signal derived from attention weights (Lu et al., 2025; Zhang et al., 2024b), encouraging attention to consistently concentrate on the trigger image. This design improves attack reliability under dynamic visual distractions.
(a) Successful case
(b) Unsuccessful case
Figure 2. Attention maps for the two cases. The red box marks the trigger image; warmer colors indicate higher attention. In the successful case, attention is concentrated on the trigger image region, whereas in the unsuccessful case, attention is dispersed across the screenshot.
We conduct an in-depth evaluation of our proposed Chameleon. We first construct six highly realistic datasets that simulate widely used websites with dynamically varying images and text. We assess four representative GUI agents (UI-TARS-7B-DPO (Qin et al., 2025), OS-Atlas-Base-7B (Wu et al., 2024b), Qwen2-VL-7B (Wang et al., 2024a), and LLaVA-1.5-13B (Liu et al., 2023)) and find that Chameleon substantially outperforms all baselines across all websites (Madry et al., 2017; Wu et al., 2024a; Aichberger et al., 2025). We further observe favorable cross-model generalization across similar models. An ablation study shows that both the LLM-Driven Environment Simulation and the Attention Black Hole are necessary for Chameleon. Finally, we evaluate several practical defenses and find that many existing strategies are ineffective against Chameleon without substantially harming user experience, highlighting the need for new defense mechanisms tailored to GUI agents.
The contributions of this paper are threefold.
  • We formalize a realistic dynamic-environment threat model for EIA. Under this novel threat model, we demonstrate that existing EIA approaches are largely ineffective.
  • We propose Chameleon, a new environmental injection attack framework that is effective under the realistic dynamic-environment threat model, enabling the discovery of previously hidden vulnerabilities in existing GUI agents.
  • We conduct extensive experiments on four representative GUI agents across six websites. The results obtained through Chameleon uncover previously underexplored vulnerabilities in existing GUI agents, highlighting the urgent need for more robust security mechanisms.
2. Background and Related Work
2.1. Large Vision-Language Models
Large Vision-Language Models are foundational to modern GUI agents. These models typically consist of three main components: a visual encoder, a connector, and a large language model. For visual input, a visual encoder, such as CLIP (Radford et al., 2021), first partitions the input image into many patches, each representing a local pixel region (e.g., forming a 14×14 grid), and then extracts corresponding visual features. Subsequently, these visual features are transformed into visual tokens via a connector module, such as a Multi-Layer Perception (MLP) or Q-Former (Li et al., 2023). These visual tokens can then be directly fed into the subsequent LLM. Typically, each visual token corresponds to multiple pixels within the input image. Popular LVLMs that adopt this architecture include the Qwen (Wang et al., 2024a; Bai et al., 2025; Hsieh and Wei, 2025) series and LLaVA (Liu et al., 2023; Guo et al., 2024) series among others.
2.2. GUI Agents
GUI agents have recently attracted growing interest in the software engineering community, as they offer a promising direction for automating interactions with real-world applications (Ye et al., 2025a; Qian et al., 2023; Zhang et al., 2025a). Building on the strong multimodal capabilities of LVLMs, researchers have developed GUI agents that automate complex tasks within graphical user interfaces. Unlike early approaches (Deng et al., 2023; Nakano et al., 2021) that directly fed raw HTML and human instructions into LLMs, an approach often hindered by redundant or irrelevant information, modern GUI agents (Qin et al., 2025; Wu et al., 2024b; Cheng et al., 2024; Hong et al., 2024) leverage LVLMs to process rendered webpage screenshots as input, achieving significantly better performance.
Recent advances further improve GUI agents by introducing techniques such as Set-of-Marks (SoMs) (Yang et al., 2023) to enhance interaction with GUI elements. Open-source models have also contributed to the field by increasing agent capabilities and reducing deployment costs. For example, OS-ATLAS (Wu et al., 2024b) utilizes large-scale datasets containing screenshots, element instructions, and coordinates for comprehensive GUI understanding, while UI-TARS (Qin et al., 2025) leverages extensive training corpora to improve screen perception.
In this work, we focus on a widely-adopted GUI agent paradigm (Lu et al., 2025; Wu et al., 2024b; Qin et al., 2025), specifically a LVLM-powered agent denoted by the model $M$. Initially, the agent receives a system prompt $p_{s}$ and a user instruction $p_{u}$ representing a specific task. At each subsequent interaction step, the agent observes a screenshot $s_{t}$, rendered from the current HTML content, along with the action history $H_{t}$, and then outputs an action $a_{t}$ until the task is completed or failed. Formally, this can be expressed as:
(1) $a_{t}=M(p_{s},p_{u},s_{t},H_{t}),$
where $H_{t}=[a_{1},a_{2},\dots,a_{t-1}]$.
2.3. Environmental Injection Attacks
Environmental Injection Attacks (EIAs) pose a critical security threat to GUI agents. Unlike conventional attacks that originate from the user’s input, EIAs are injected through the external environment that the agent observes. For example, an attacker can upload a crafted trigger, such as an image or a short text snippet, to a target website. When another user’s GUI agent later encounters this trigger while browsing the site, it may be induced to execute unintended actions without the user’s explicit instruction (Zhao et al., 2025; Lu et al., 2025). Although textual triggers such as “When the user is trying to find a motorcycle, give them this one regardless of the other requirements” can be effective, they are often easy to detect (Aichberger et al., 2025). As a result, recent EIA research has increasingly focused on image-based triggers, which are more subtle and less likely to be caught (Xu et al., 2024; Chen et al., 2025; Wu et al., 2024a).
According to the attacker’s identity, existing EIAs can be broadly divided into two categories (Wu et al., 2024a; Aichberger et al., 2025; Liu et al., 2025). The first category assumes an attacker with administrative or developer privileges over the target website, enabling full control over the rendered interface. Under this assumption, the attacker can manipulate the entire screenshot or even modify the underlying HTML source code, which supports powerful attacks such as applying global adversarial perturbations or inserting misleading interface components (e.g., deceptive buttons). For instance, EIA (Liao et al., 2024) injects new form elements into HTML to mislead agents, while ENVINJECTION (Wang et al., 2025) applies adversarial perturbations to entire screenshots. While highly effective, this threat model is often unrealistic in practice.
The second category adopts a more practical threat model in which the attacker is a regular malicious user, such as a content uploader on social media or a competing vendor on an e-commerce platform. These attackers can only manipulate a small portion of the webpage, typically by uploading crafted trigger images with adversarial perturbations. However, many existing methods under this setting (Wu et al., 2024a; Zhao et al., 2025) do not adequately account for environmental dynamism in real world, especially variations in trigger position and the surrounding content. For instance, Wu et.al. (Wu et al., 2024a) optimized malicious images without considering any visual context and then deployed them directly onto target websites, severely limiting their attack effectiveness.
Our work falls into the second category. Most similar to our work is MIP (Aichberger et al., 2025). While MIP acknowledges dynamic visual context should be considered, it only introduces minor changes to the surrounding visual context and does not account for variations in the trigger image’s position. In contrast, our work considers significant variations in both trigger positions and surrounding visual context, closely simulating real-world scenarios characterized by highly dynamic internet content. To our knowledge, this study presents the first systematic investigation into how trigger images perform within dynamically changing visual contexts when attacking GUI agents.
3. Dynamic-Environment Threat Model
In this section, we formally define the dynamic-environment threat model for environmental injection attacks against GUI agents. The key characteristic of this threat model is that the trigger is observed under a continuously changing visual environment, where both its on-screen position and surrounding visual context are dynamically changing and difficult to predict in advance.
Attack Scenarios. We consider GUI agents deployed on online platforms such as social media services and e-commerce websites, where users can upload images and browse content posted by others. In these platforms, the rendered pages seen by GUI agents are not static. The displayed content can change over time due to common platform behaviors, including updates to rankings and recommendations, advertisement placement, and frequent content refresh. Consequently, a trigger uploaded by an attacker may be displayed at different locations and within different surrounding contexts when viewed by different users or at different times.
Attacker’s Goal. The attacker aims to craft a trigger image and upload it to the target website. When other users later interact with the website through GUI agents, the trigger is rendered as part of the observed screenshot and is intended to steer the agent toward attacker-chosen behavior, such as clicking a promotional link, visiting a malicious URL, or selecting a specific item. Formally, the goal is to induce the agent to output an incorrect target action when the trigger is present.
Attacker’s Constraints. The attacker has no administrative privileges and cannot modify the website source code or the victim’s device. More importantly, the attacker cannot control or reliably predict the trigger’s on-screen position or its surrounding visual context, both of which may differ across users and over time. Additionally, attackers have no knowledge of the GUI agent’s action history or the specific user instructions provided to the agent.
Attacker’s Capabilities. Attackers can upload trigger images to a target website. Because website layouts and styles typically remain stable over short periods, attackers can reliably anticipate the overall structure and appearance of the target page. Furthermore, following prior work (Aichberger et al., 2025; Wu et al., 2024a), we assume that attackers have white-box access to the model powering the GUI agent  [^2] We also consider transfer-based black-box attacks in our evaluation., including knowledge of its gradients and architecture. We consider this assumption realistic because many GUI agents are deployed locally using open-weight LVLMs by users, allowing attackers to directly craft triggers against these publicly available models. Given the increasing deployment of GUI agents and the ease with which trigger images can be distributed online (Wang et al., 2024b; Tur et al., 2025), we think compromising even one popular model may lead to far-reaching consequences.
4. Methodology
Under the realistic dynamic-environment threat model, our results in Section 5.2 show that existing environmental injection attacks have very low effectiveness, and therefore fail to fully expose the vulnerabilities of GUI agents in real-world settings.
To tackle the challenge posed by dynamic environments, we propose Chameleon, a novel attack framework for LVLM-powered GUI agents. We begin by describing the overall attack pipeline in Section 4.1. We then present LLM-Driven Environment Simulation in Section 4.2, which automatically generates large-scale realistic training data that captures dynamic webpage variations. Finally, we introduce Attention Black Hole in Section 4.3, which ensure more effective training by encouraging the model to consistently focus its attention on the trigger region. An overview of Chameleon is shown in Figure 3.
Figure 3. Overview of our proposed Chameleon.
4.1. Overview
For a chosen target website, such as a social media platform like RedNote  [^3] https://www.xiaohongshu.com/ or an e-commerce site like Amazon  [^4] https://www.amazon.com/, the attacker first selects a benign-looking trigger image $I$, such as a product photo or a post cover. This image serves as the starting point and will be optimized during training.
At each training step, we first use our LLM-Driven Environment Simulation $G_{\mathrm{LES}}$ (Section 4.2) to construct training data. Specifically, it generates a webpage screenshot $s$ that contains the perturbed trigger image, a corresponding user instruction $p_{u}$, and a pixel-level mask $mask^{shot}$ that identifies the trigger region within $s$. We then update the perturbation $\delta$ inside the trigger region using Projected Gradient Descent (PGD) (Madry et al., 2017). To ensure the resulting malicious trigger image remains imperceptible to users, we constrain the perturbation $\delta$ to a predefined $\ell_{\infty}$ norm bound $\epsilon$, formally:
(2) $|\delta|_{\infty}\leq\epsilon.$
[…]
(3) $\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{CE}}(a,\hat{a})+\lambda\cdot\mathcal{L}_{\mathrm{attn}},$
where $\lambda$ controls the relative weight of the attention loss.
After the training procedure described above, the attacker uploads the resulting trigger image to the target website, acting as a regular user. Once the trigger image is integrated into a public post or product listing, any GUI agent browsing the site and encountering the trigger image within a webpage will be susceptible to attack, potentially performing the intended harmful action. Algorithm 1 summarizes the overall procedure of Chameleon, covering the training and deployment stages.
Input: Target website URL, raw trigger image $I$, GUI agent $M$, system prompt $p_{s}$, target action $\hat{a}$, perturbation bound $\epsilon$, step size $\alpha$, weight $\lambda$, training steps $K$, flagship LLM
/* Training */
$T\leftarrow$ Parse target website with flagship LLM;
/* HTML template */
$C\leftarrow$ Crawl large-scale multimodal content;
/* image-text corpus */
1 Initialize perturbation $\delta\leftarrow 0$ ;
2 for $k=1$ to $K$ do
  /* LLM-Driven Environment Simulation (LES) */
3  $(s,mask,p_{u})\leftarrow G_{LES}(T,C,I+\delta)$ ;
4  $(a,A)\leftarrow M(p_{s},p_{u},s)$ ;
  /* Attention Black Hole (ABH) */
5  $\mathcal{L}_{attn}\leftarrow\textsc{ABH}(A,mask)$ ;
[…]
7  $\mathcal{L}\leftarrow\mathcal{L}_{CE}+\lambda\mathcal{L}_{attn}$ ;
8  $\delta\leftarrow\text{Update}(\delta,\mathcal{L},\epsilon,\alpha)$ ;
9
10 end for
/* Deployment */
11 $I_{adv}\leftarrow I+\delta$ ;
12 Upload $I_{adv}$ to target website ;
Algorithm 1 Chameleon
4.2. LLM-Driven Environment Simulation
We argue that enabling the trigger to generalize across dynamic environments requires training over a large collection of realistic visual environments. To address this challenge, we introduce LLM-Driven Environment Simulation (LES). LES leverages the generative capabilities of LLMs to automatically synthesize a large number of realistic training samples. By systematically varying the trigger’s position and its surrounding visual context during synthesis, LES ensures the optimized trigger learns to remain effective under continuously changing environments. Concretely, we define a generation function $G_{LES}$ that outputs a rendered webpage screenshot, a binary mask indicating the trigger location, and a user instruction for downstream training.
First, to ensure that our simulated environments are realistic, we construct a high-fidelity HTML template $T$ for each target website. We begin by collecting the raw HTML from the live website and treat it as the initial reference for both layout and styling. Our goal is to remove concrete content (e.g., product titles, prices, user posts, and recommendations) while preserving the core structural and stylistic components (e.g., DOM hierarchy, CSS classes, layout containers, and visual themes), so that the resulting template remains visually indistinguishable from the real site after being populated. To achieve this, we leverage a flagship LLM in a multi-turn, interactive refinement process. In each round, we provide the LLM with the raw HTML and instruct it to abstract away concrete content into clearly defined placeholders that can later be filled with arbitrary image-text pairs. We then render the produced template, inspect whether any real content is still hard-coded or whether the layout has been unintentionally altered, and feed this feedback back to the LLM for further refinement. After several rounds, this process yields a clean, fillable HTML template $T$ that preserves the authentic look of the target website while allowing us to programmatically inject diverse content during simulation. Notably, for each target website, this template construction is a one-time effort, and a single HTML template $T$ can be reused to generate all simulated environments for that website.
We then construct a large-scale, multimodal corpus $C$ by crawling the target website for actual content, gathering an average of over 5,000 distinct items (e.g., product images, titles, prices) for each website. This process ensures our content pool is not only realistic but also sufficiently diverse to simulate a wide range of scenarios.
Next, we define an HTML generation function $g$, which integrates randomly selected contextual image-text pairs from corpus $C$ along with the pre-selected trigger image $I$ into the HTML template $T$, producing renderable HTML code $h$:
(4) $h=g(T,C,I).$
The generated HTML code is subsequently rendered into a screenshot $s$ accompanied by a binary pixel-level mask $mask^{shot}$ marking the trigger region, where pixels corresponding to the trigger image region are set to 1, and all remaining pixels are set to 0:
(5) $s,mask^{shot}=f(h).$
Considering the inherently unpredictable nature of real user instructions during actual usage, we utilize another advanced LLM to automatically generate realistic and diverse user instructions $p_{u}$ corresponding to each synthesized screenshot $s$. Thus, the complete environment generation process within our LES framework can be formally described as:
(6) $s,mask^{shot},p_{u}=G_{LES}(f(g(T,C,I))).$
Through this comprehensive procedure, LES effectively combines known webpage structures with realistic and diverse environment generation, thereby significantly enhancing the generalizability of optimized trigger images under dynamically varying web environments.
4.3. Attention Black Hole
After obtaining large-scale realistic training data, the next challenge is how to optimize the trigger more effectively. We argue that reliable attacks in dynamically changing visual environment require the model to consistently prioritize the trigger over competing interface elements. Thus, we propose Attention Black Hole (ABH). By converting attention weights into explicit supervisory signals, ABH encourages the agent to consistently focus on the trigger region, thereby improving robustness by reducing interference from distracting content in the dynamic environment.
Suppose that after passing through the visual encoder and connector, the input screenshot is transformed into a sequence of image tokens with length $n\times m$. Given the text tokens comprising system prompt $p_{s}$, user instruction $p_{u}$, and action history $H_{t}$ along with image tokens, the LVLM generates a new token sequence $N$. We utilize attention weights from the last layer of the LVLM to quantify the interactions between each newly generated token and all image tokens, producing an attention map $A_{i,j}$ defined as follows:
(7) $A_{i,j}=\frac{1}{|N|\times H}\sum_{t=1}^{|N|}\sum_{h=1}^{H}Attention_{t,h}^{i,j},$
where $Attention_{t,h}^{i,j}$ denotes the attention weight from the $h$-th attention head, relating the $t$-th new token to the $(j+m\times(i-1))$-th image token. $H$ represents the number of attention heads, and $|N|$ indicates the total number of newly generated tokens.
Based on the binary mask $mask^{shot}$ of the trigger image at the pixel level, we apply a resizing operation to obtain a token-level binary mask $mask^{attn}$ of size $n\times m$, where $mask^{attn}_{i,j}$ is set to 1 if the $(j+m\times(i-1))$-th image token corresponds to a patch that overlaps with the trigger image region, and 0 otherwise.
Subsequently, we define $\overline{A}_{in}$, representing the model’s degree of focus on the trigger image region, as the mean attention weight within the trigger image region. Likewise, $\overline{A}_{out}$ represents the model’s attention to other regions of the screenshot and is computed as the mean attention weight outside the trigger region. Formally, these are given by:
(8) $\overline{A}_{in}=\frac{\sum_{i,j}A_{i,j}\times mask^{attn}_{i,j}}{\sum_{i,j}mask^{attn}_{i,j}},$
(9) $\overline{A}_{out}=\frac{\sum_{i,j}A_{i,j}\times(1-mask^{attn}_{i,j})}{\sum_{i,j}(1-mask^{attn}_{i,j})}.$
Finally, we define the loss function $\mathcal{L}_{attn}$ as the ratio of the average attention outside of the trigger region to the average attention within it. Formally, $\mathcal{L}_{attn}$ is defined as:
(10) $\mathcal{L}_{attn}=\frac{\overline{A}_{out}}{\overline{A}_{in}}.$
Minimizing this loss explicitly guides the model to focus on the trigger region, which is crucial for disregarding the distractions caused by the dynamic environment and maintaining attack robustness.
5. Experiments
To systematically evaluate our proposed Chameleon, we conduct extensive experiments designed to answer the following Research Questions (RQs).
Figure 4. Illustrative screenshots for each website. Trigger images are outlined in red.
First, as stated in our proposed threat model, the primary goal of an attacker is to mislead the GUI agent into executing an incorrect action when exposed to the trigger image. We therefore design our first research question to evaluate the fundamental effectiveness of Chameleon in achieving this objective.
RQ1: How effective is Chameleon for conducting environmental injection attacks against GUI agents? To answer this, we evaluate the attack success rates of Chameleon on six distinct and well-known target websites.
Although compromising even a single popular model can pose a significant security risk due to the vast user base of the internet, an attacker’s impact further increases if a trigger crafted against one LVLM transfers to unseen LVLMs. We therefore focus exclusively on cross-model transferability.
RQ2: How transferable is Chameleon to unseen LVLMs? We train triggers on a surrogate LVLM and evaluate them—without any adaptation—on multiple target LVLMs to measure cross-model transferability.
Thirdly, we aim to understand the individual contributions of our core technical innovations. We therefore design our third research question to analyze the importance of each component within our framework.
RQ3: How do the LLM-Driven Environment Simulation and Attention Black Hole contribute to the performance of Chameleon? We perform an ablation study to isolate and quantify the impact of each of these components on the overall effectiveness of the attack.
Finally, we explore whether existing defenses can mitigate Chameleon under our realistic dynamic-environment threat model, and what practical mitigation strategies may be effective. We therefore design our fourth research question to study potential defenses.
RQ4: What defenses are effective against Chameleon under the dynamic-environment threat model? We evaluate a set of practical defense strategies and measure their impact on attack success rate, aiming to identify promising directions for securing GUI agents in realistic deployments.
5.1. Experimental Setup
Datasets. To ensure the realism of our evaluation, we selected six target websites spanning three representative categories of GUI-based online services, as summarized in Table 1. These categories reflect diverse user interaction patterns, which are critical for assessing the effectiveness of Chameleon across practical use cases. For each website, we construct a validation set of 300 and a test set of 1,200 screenshot-instruction pairs using the LLM-Driven Environment Simulation introduced in Section 4.2. Importantly, the images and instructions in the training, validation, and test sets are mutually disjoint, ensuring rigorous evaluation. Figure 4 displays some example screenshots for each of the six websites.
Table 1. Datasets used in this work.
Category Website URL
Shopping Amazon https://www.amazon.com/
Taobao https://www.taobao.com/
Social Media RedNote https://www.xiaohongshu.com/
Bilibili https://www.bilibili.com/
Music Streaming NetEase Cloud Music http://ir.music.163.com/en/
QQ Music https://y.qq.com/
LVLMs for GUI agents. Our evaluation is conducted on a diverse set of four popular and representative Large Vision-Language Models. These include two models specialized for GUI tasks, UI-TARS-7B-DPO (Qin et al., 2025) and OS-Atlas-Base-7B (Wu et al., 2024b), as well as two general-purpose LVLMs, Qwen2-VL-7B (Wang et al., 2024a) and LLaVA-1.5-13B (Liu et al., 2023). Notably, UI-TARS-7B-DPO and OS-Atlas-Base-7B are fine-tuned from Qwen2-VL-7B.
Baselines. Although a large body of prior work (Lu et al., 2025; Zhao et al., 2025; Wang et al., 2025; Aichberger et al., 2025) studies environmental injection attacks, we find that most of them cannot be directly applied under our dynamic-environment threat model. Many approaches assume an attacker with administrative privileges that are not practical, while others rely on explicit textual injections that are easily filtered. After carefully filtering the literature to match our attacker capabilities and deployment constraints, we select two representative baselines:
  • PGD (Wu et al., 2024a). We adopt a standard PGD-based adversarial attack (Madry et al., 2017) as a baseline, following prior trigger-image EIA studies such as (Wu et al., 2024a). During trigger optimization, this baseline does not incorporate any environmental context, including surrounding webpage content or stylistic information.
  • MIP (Aichberger et al., 2025). We implement MIP, which constructs training samples by manually creating dynamic environments and then optimizes the trigger using PGD. Following its experimental setup, for each trigger image we manually craft 12 screenshots with varying contexts for optimization.
Target Action. Consistent with the attacker’s goal, we define the target action as instructing the agent to navigate to a specific malicious URL. This could be a URL for a promotional site or a more harmful phishing page. For each target website, we set the target action to a navigation command that directs the agent to a malicious URL such as:
goto [http://one-example.com]
Evaluation Metrics. We employ the Attack Success Rate (ASR) to evaluate the effectiveness of Chameleon, formally defined as follows:
(11) $\text{ASR}=\frac{N_{\text{attack}}}{N_{\text{total}}},$
where $N_{\text{attack}}$ denotes the number of responses matching the target action within the test set, and $N_{\text{total}}$ represents the total size of the test set. A strict string matching criterion is adopted to verify whether the agent’s response matches the target action precisely.
Implementation Details. We employ GPT-4o (Achiam et al., 2023) to parse the live website, preserving its core structural and stylistic components while removing existing content, and we use Qwen2.5-VL-32B-Instruct (Bai et al., 2025) to automatically generate realistic and diverse user instructions. The system prompt and user prompt for the GUI agent, as well as the agent’s action space, are adapted from VisualWebArena (Koh et al., 2024). Since it is challenging to collect authentic action histories, we follow ENVINJECTION (Wang et al., 2025) and randomly sample 0 to 10 historical actions for each instance. To ensure rigorous evaluation, we maintain a strict separation, with no overlap in the sampled action histories used across the training, validation, and testing. The hyperparameter $\lambda$ is set to $0.3$ and $\epsilon$ is set to $\frac{32}{255}$. The perturbation is optimized for a total of $5,000$ steps, with updates performed at each step using a fixed step size of $\alpha=\frac{1}{255}$.
Table 2. ASRs (%) of baselines and Chameleon across different GUI agents and websites. NetEase is used as an abbreviation for NetEase Cloud Music. Gain denotes the improvement of Chameleon over the best-performing baseline for each setting.
Model Method Shopping Social Media Music Streaming Avg.
Amazon Taobao RedNote Bilibili NetEase QQ Music
UI-TARS-7B-DPO PGD 3.17 2.83 4.08 5.75 6.58 1.25 3.94
MIP 5.67 4.42 5.17 7.17 11.08 1.33 5.78
Chameleon 20.75 22.58 23.25 41.58 41.08 8.08 26.22
Gain $\times$3.66 $\times$5.11 $\times$4.50 $\times$5.80 $\times$3.71 $\times$6.08 $\times$4.54
OS-Atlas-Base-7B PGD 4.25 5.17 3.83 7.08 8.00 3.25 5.26
MIP 6.17 7.25 3.92 9.92 10.08 3.50 6.81
Chameleon 23.67 33.17 26.42 35.50 43.75 33.08 32.60
Gain $\times$3.84 $\times$4.58 $\times$6.74 $\times$3.58 $\times$4.34 $\times$9.45 $\times$4.79
Qwen2-VL-7B PGD 2.50 4.00 4.67 6.83 7.58 2.08 4.61
MIP 6.75 8.17 7.08 11.50 14.17 4.17 8.64
Chameleon 9.42 14.42 17.58 22.42 25.93 9.92 16.62
Gain $\times$1.40 $\times$1.76 $\times$2.48 $\times$1.95 $\times$1.83 $\times$2.38 $\times$1.92
LLaVA-1.5-13B PGD 4.75 6.25 7.08 8.17 8.75 3.50 6.42
MIP 9.17 10.17 7.92 14.67 15.25 9.42 11.1
Chameleon 37.17 43.00 60.75 50.83 76.58 32.08 50.07
Gain $\times$4.05 $\times$4.23 $\times$7.67 $\times$3.46 $\times$5.02 $\times$3.41 $\times$4.51
5.2. RQ1: Effectiveness in Attacking Target GUI Agents
In this RQ, we evaluate whether Chameleon can reliably mislead GUI agents into executing a predefined malicious action when an adversarial trigger is embedded in a target website under the dynamic-environment threat model.
Setup. We compare the ASR of Chameleon with two baselines across four representative GUI agents and six target websites described in our experimental setup.
Results. Table 2 reports the ASRs (%) of all methods across models and websites. Higher ASR indicates stronger attack effectiveness.
Analyses.Existing attacks fail to fully expose GUI agent vulnerabilities under realistic dynamic environments. Although prior studies reported relatively high ASRs for baseline attacks under static or slightly varying settings (Wu et al., 2024a; Aichberger et al., 2025), Table 2 shows that both PGD and MIP remain largely ineffective in our dynamic-environment evaluation, with ASRs remaining close to zero across models and websites. For example, on OS-Atlas-Base-7B, PGD achieves only 5.26% ASR on average and MIP reaches 6.81%, suggesting that these approaches substantially underestimate the vulnerability of GUI agents when evaluated in realistic, continuously changing web contexts. ❷ Chameleon achieves substantially higher attack effectiveness across all configurations. In contrast to the baselines, Chameleon consistently improves ASR by a large margin for every evaluated agent and website. For instance, on OS-Atlas-Base-7B, Chameleon increases the average ASR from 5.26% (PGD) and 6.81% (MIP) to 32.60%. Similar gains are observed on UI-TARS-7B-DPO (from 3.94% and 5.78% to 26.22%) and Qwen2-VL-7B (from 4.61% and 8.64% to 16.62%). These results demonstrate that Chameleon can effectively uncover previously hidden vulnerabilities of GUI agents that remain unexposed by existing attacks under realistic dynamic environments.
Answer to RQ1: Under the dynamic-environment threat model, existing attacks exhibit very low effectiveness and thus fail to fully reveal GUI agent vulnerabilities, whereas Chameleon achieves substantially higher ASRs across all models and websites.
5.3. RQ2: Generalization Ability Across LVLMs
In this RQ, we evaluate the generalization ability of our trigger image across multiple LVLMs. Specifically, we aim to examine whether a trigger trained on one LVLM-powered GUI agent can generalize to other unseen agents in a black-box setting.
(a) NetEase Cloud Music
(b) Amazon
(c) RedNote
Figure 5. Transferability of Chameleon across models. Each cell represents the ASR (%) when the trigger image is trained on the surrogate model (row) and tested on the target model (column).
Setup. We conduct a $4\times 4$ transfer experiment across four LVLMs: UI-TARS-7B-DPO, OS-Atlas-Base-7B, Qwen2-VL-7B and LLaVA-1.5-13B. For each surrogate model, we optimize the trigger with white-box access to that model only, then evaluate it zero-shot on the remaining three target models under black-box access. Transferability is quantified by the ASR.
Results. The cross-model transferability results are shown in Figure 5. Due to space limitations, we present results only on three representative target websites, namely NetEase Cloud Music, Amazon, and RedNote.
Analyses.Transfer is stronger between related models. OS-Atlas-Base-7B and UI-TARS-7B-DPO are both fine-tuned from Qwen2-VL-7B, and we observe moderate bidirectional transfer within this family. For instance, Qwen2-VL-7B $\to$ OS-Atlas-Base-7B achieves 14.11% and OS-Atlas-Base-7B $\to$ Qwen2-VL-7B achieves 19.17%. These results indicate that shared architectures and training data lead to overlapping representations, which can be exploited by adversarial triggers. This observation underscores a practical security concern: a trigger optimized against one model can remain effective against other closely related variants, thereby broadening the potential impact of our proposed Chameleon. ❷ Transfer collapses across dissimilar models. In contrast, LLaVA-1.5-13B differs substantially from the Qwen2-VL family in both architecture and training data. As a result, all transfer pairs involving LLaVA-1.5-13B yield 0.00% ASR in our experiments. This sharp contrast highlights the limits of transferability when model families differ substantially, suggesting that architectural heterogeneity can serve as a natural barrier against cross-model attacks, although it does not eliminate risks in the widely used internet ecosystem.
Answer to RQ2: Chameleon transfers well between similar models but shows negligible transfer across dissimilar ones (e.g., involving LLaVA-1.5-13B).
5.4. RQ3: Ablation Study
In this RQ, we investigate the contribution of the LLM-Driven Environment Simulation (LES) and the Attention Black Hole (ABH) to the effectiveness of Chameleon.
Table 3. ASRs (%) under different ablation settings. NetEase is used as an abbreviation for NetEase Cloud Music.
Method Shopping Social Media Music Streaming Avg.
Amazon Taobao RedNote Bilibili NetEase QQ Music
OS-Atlas-Base-7B
Chameleon 23.67 33.17 26.42 35.50 43.75 33.08 32.60
w/o LES 14.89 (-8.78) 15.26 (-17.91) 13.28 (-13.14) 20.81 (-14.69) 20.11 (-23.64) 15.73 (-17.35) 16.68 (-15.92)
w/o ABH 17.97 (-5.7) 29.59 (-3.58) 24.10 (-2.32) 30.02 (-5.48) 39.77 (-3.98) 26.76 (-6.32) 28.20 (-4.4)
LLaVA-1.5-13B
Chameleon 37.17 43.00 60.75 50.83 76.58 32.08 50.07
w/o LES 20.75 (-16.42) 29.75 (-13.25) 34.42 (-26.33) 35.67 (-15.16) 60.50 (-16.08) 23.09 (-8.99) 34.03 (-16.04)
w/o ABH 26.83 (-10.34) 34.92 (-8.08) 53.50 (-7.25) 46.71 (-4.12) 73.00 (-3.58) 20.75 (-11.33) 42.62 (-7.45)
Setup. We conduct an ablation study on OS-Atlas-Base-7B and LLaVA-1.5-13B by evaluating three settings across six websites: ❶ Chameleon: the full approach with both LES and ABH; ❷ Removing LES: the trigger image is trained on a limited set of 100 manually collected screenshots, without the automatic and scalable context construction provided by LES; ❸ Removing ABH: the trigger image is trained with only the cross-entropy loss $\mathcal{L}_{\mathrm{CE}}$, omitting $\mathcal{L}_{\mathrm{attn}}$.
Results. The detailed results for ablation study are presented in Table 3.
Analyses.Contribution of LES. Removing LES leads to a consistent reduction in ASR across all websites. For example, with OS-Atlas-Base-7B on NetEase Cloud Music, the ASR decreases from 43.75% to 20.11% when LES is removed. Similarly, with LLaVA-1.5-13B on RedNote, the ASR drops from 60.75% to 34.42%. These results indicate that large-scale automatic simulation of realistic and diverse visual contexts during training is critical for achieving effective attacks. ❷ Contribution of ABH. Removing ABH also reduces ASR on every website, demonstrating that explicit attention supervision is consistently beneficial. For instance, with OS-Atlas-Base-7B on QQ Music, the ASR decreases by 6.32%. For LLaVA-1.5-13B, removing ABH reduces ASR by 10.34% on Amazon and by 4.12% on Bilibili. We hypothesize that ABH improves robustness by suppressing interference from non-trigger regions and encouraging the model to maintain stable focus on the trigger across dynamically changing environments.
Answer to RQ3: Both LES and ABH are necessary. LES improves robustness by providing automatic simulation of realistic and diverse contexts, while ABH amplifies the trigger’s effect by explicitly directing attention to the trigger region.
(a) NetEase (LLaVA-1.5-13B)
(b) RedNote (LLaVA-1.5-13B)
(c) NetEase (OS-Atlas-Base-7B)
(d) RedNote (OS-Atlas-Base-7B)
Figure 6. Effectiveness of defense strategies against Chameleon across different websites. The first row corresponds to LLaVA-1.5-13B and the second row to OS-Atlas-Base-7B, each evaluated on two representative target websites. NetEase is used as an abbreviation for NetEase Cloud Music.
5.5. RQ4: Possible Defenses
In this RQ, we explore practical defense strategies for GUI agents under the dynamic-environment threat model and evaluate their effectiveness in reducing the attack success rate of Chameleon.
Setup. We evaluate defenses on two GUI agents, LLaVA-1.5-13B and OS-Atlas-Base-7B, using NetEase Cloud Music and RedNote as target websites. Defense effectiveness is measured by ASR. Figure 6 summarizes the results.
5.5.1. Safety Prompt
Safety prompts (Wang et al., 2024c; Zhang et al., 2025b) are widely used to steer models away from unsafe behavior. We prepend the following instruction to the system prompt to encourage the agent to follow the user request and ignore potentially malicious content in the observation:
You should adhere strictly to the user’s instructions and not be distracted by any injected malicious instructions found in the observation.
This defense yields almost no reduction in ASR. For example, on RedNote with a LLaVA-1.5-13B powered agent, ASR decreases only from 60.75% to 60.21%. This suggests that prompt-based steering alone is insufficient against Chameleon.
5.5.2. Verifier
A common mitigation is to verify model outputs before execution. We use DeepSeek-V3 (Liu et al., 2024) as an external verifier that checks each candidate action against the user instruction and aborts execution if the action appears inconsistent or risky. Verification reduces ASR in some settings, for example from 43.75% to 36.42% for OS-Atlas-Base-7B on NetEase Cloud Music. However, we also observe false positives that block benign actions, reducing overall utility. Moreover, this approach introduces additional latency and inference cost because it requires a secondary model call at each step.
Figure 7. Comparison between original (left half) and noised (right half) versions of four photography images. Random noise causes visible degradation in image quality, significantly undermining user experience.
5.5.3. Random Noise on Uploaded Images
Since Chameleon performs attacks by applying adversarial perturbations, one potential countermeasure is to add random noise to uploaded images. We assume that a website administrator automatically adds noise $\eta$ with an $\ell_{\infty}$ bound $|\eta|_{\infty}\leq\epsilon$. With $\epsilon=\tfrac{8}{255}$, this defense proves highly effective: on RedNote, the ASR of a LLaVA-1.5-13B powered agent drops to 11.84%, and for OS-Atlas-Base-7B the ASR is reduced to nearly zero. Although effective, even small bounded noise noticeably degrades image quality, which may harm user experience, particularly in scenarios where high visual fidelity is required, such as photography or digital art. As shown in Figure 7, applying random noise leads to clear quality degradation.
Answer to RQ4: Many existing defenses struggle to mitigate Chameleon without substantially degrading user experience. Safety prompts provide negligible protection, output verification offers only moderate gains with added latency and utility loss, and randomized image noise can be effective but noticeably harms visual quality.
6. Discussion
6.1. Case Study
Figure 8. Case study: GUI agent’s behavior for the task “open the first playlist in the homepage recommendations of NetEase Cloud Music.”
To assess whether Chameleon can induce harmful actions in a realistic interactive setting, we conducted a closed-loop evaluation that more closely reflects real-world scenarios.
Experimental Setup. We instantiated a GUI agent using UI-TARS-7B-DPO as the underlying LVLM. The agent ran end-to-end without human assistance, issuing actions that directly controlled a browser. To avoid any risk to production services, all experiments were performed in a fully isolated sandbox. Specifically, we constructed a static website that mirrors the visual layout and interaction patterns of NetEase Cloud Music, deployed on localhost inside a containerized network where all outbound traffic was blocked at the firewall and DNS resolution was disabled. The agent’s initial task was: “open the first playlist in the homepage recommendations of NetEase Cloud Music.” The agent launched the browser, entered the public NetEase Cloud Music URL as a string, and the request was internally rewritten to the local mirror, which returned the sandboxed homepage. We instrumented the environment to record screenshots and the agent’s actions at every step.
The homepage contained a single trigger image embedded in the recommendation rail. We evaluated two conditions that were identical in layout and ordering of items. In the without attack condition, the trigger image was the original clean image. In the with attack condition, the same image carried the perturbation optimized by Chameleon under an $\ell_{\infty}$ bound $\epsilon$. The agent observed the rendered page’s screenshot, action history, and the browser state updated accordingly, which completed the control loop. Figure 8 shows the sandboxed homepage in the bottom-left panel, the target playlist page in the top-right panel, and the promotional site used as the malicious target in the bottom-right panel. Success for the benign task was defined as navigating to the first playlist page. Attack success was defined as issuing a navigation action that opened the predefined promotional webpage.
Results and Analyses. As shown in Figure 8, in the without attack condition, the agent executed the instructed task correctly. It parsed the homepage’s screenshot, selected the first recommended playlist, and navigated to the corresponding playlist detail page. In the with attack condition, the agent encountered the homepage containing the perturbed trigger image and immediately issued a navigation action that opened the promotional site, rather than proceeding to the first playlist. These observations demonstrate that Chameleon is not restricted to open-loop or purely simulated pipelines. The attack transfers to a closed-loop setting where actions modify state and state feeds back to the policy. The outcomes highlight tangible risks to Internet-facing GUI agents, since a single user-uploaded image can redirect goal-directed behavior even when the surrounding interface and task specification remain benign.
6.2. Threats to Validity
We identify four primary threats to the validity of our study:
Evaluation Metrics. We compute ASR using a strict string-matching criterion that counts an attack as successful only when the exact target action is produced. Minor variations such as additional spaces, a www prefix, or appended paths are treated as failures. This conservative definition provides a lower bound on practical risk. Nevertheless, even with this lower-bound measure, our results reveal widespread vulnerabilities in GUI agents, suggesting that this limitation does not undermine our conclusions. To further mitigate this concern, future evaluations could incorporate human judgment or LLM-based assessment of semantic equivalence, which would capture cases where the produced actions are functionally equivalent despite surface-level variations.
Datasets. Existing GUI agent datasets often present webpages where text and images are static, which misaligns with real-world internet scenarios. To address this, we construct test sets using our LLM-Driven Environment Simulation, which enables large-scale generation of realistic and diverse visual contexts. Moreover, to eliminate potential data leakage, we ensure that the screenshots used in the training, validation, and test sets do not share any uploaded images. This separation guarantees that performance improvements cannot be attributed to memorization of specific samples.
Replication of Our Experiments. The behavior of LVLM-powered agents can be influenced by multiple factors, such as decoding temperature, which complicates replication. To support reproducibility, we release detailed descriptions of our experimental settings, including hyperparameters and environment configurations, along with links to the exact model checkpoints used. In addition, our full code repository is publicly available to facilitate independent verification. These measures collectively enhance transparency and provide a reliable foundation for reproducing our findings. Future work may also consider standardized benchmarks and controlled evaluation environments, which would further reduce randomness and strengthen replicability.
Ethics Statement. In this work, we propose a more practical and realistic threat model and an effective attack approach. However, our goal is not to promote malicious behavior, but rather to reveal the hidden vulnerabilities of widely used GUI agents when deployed in real-world, open-ended internet environments. By exposing these vulnerabilities, we aim to raise awareness of the potential risks and emphasize the urgent need for robust and practical defenses. We hope that our findings will inform future research on building safer and more trustworthy web-based agent systems.
7. Conclusion
In this paper, we present a more realistic threat model in which the attacker is a regular user who can only upload trigger images that appear within a dynamically changing environment. To address the challenges posed by dynamic environment, we propose Chameleon, a novel attack framework that introduces two key novelties:LLM-Driven Environment Simulation, which enables large-scale automatic generation of realistic and diverse webpage simulations, and Attention Black Hole, which explicitly guides the agent’s focus toward the trigger region. Extensive experiments across multiple websites and models demonstrate that Chameleon significantly outperforms existing methods in attack success rate. Ablation and closed-loop evaluations further confirm the effectiveness and real-world applicability of the proposed techniques. Through an evaluation of several commonly used defense strategies, we find that existing defense strategies fail to effectively mitigate the threat of Chameleon without compromising the utility of the agent. Overall, our study reveals the inherent vulnerabilities in widely used LVLM-powered GUI agents. Future work may explore automated trigger detection and effective defenses that preserve usability in open-world web environments.
8. Data Availability
The source code and access details for the datasets can be accessed at: https://github.com/zhangyitonggg/attack2gui.